Calculate the wavelength of light emitted during the transition of an electron in b/w two levels of the Li2+ ion, whose sum is 4 but difference is 2.

Grade:

2 Answers

Sunil Kumar FP
9 years ago
let the wavelength of light be represented by ^
here the sum fo two transition n1+n2 =4 and the difference is n1-n2 =2
solving both the equation we get
2n1 =6 ,n1=3
n2=1
therefore,1/^ =RZ^2 (1/n2^2 -1/n1^2)
1/^=109677 *9(1-1/9)
1/^ =109677*9*8/9
1/^ =109677*8
^=1/109677*8
^=1.139 *10^-6 cm
^=1.139*10^-8 m
^=114 A0
ninad kumbhojkar
9 years ago
sum is 4 and diff is 2
let them be ‘x’ and ‘y’
then x+y=4
        x-y=2
solving this we get 
x=3
y=1
then simply apply rydberg’s formula
and calculate